# Extraction of nonlinearity in neural networks and model compression with Koopman operator

### v1

@@ Abstract
Nonlinearity plays a crucial role in deep neural networks. In this paper, we first investigate the degree to which the nonlinearity of the neural network is essential. For this purpose, we employ the Koopman operator, extended dynamic mode decomposition, and the tensor-train format. The results imply that restricted nonlinearity is enough for the classification of handwritten numbers. Then, we propose a model compression method for deep neural networks, which could be beneficial to handling large networks in resource-constrained environments. Leveraging the Koopman operator, the proposed method enables us to use linear algebra in the internal processing of neural networks. We numerically show that the proposed method performs comparably or better than conventional methods in highly compressed model settings for the handwritten number recognition task.
@@ 1 Introduction
As is well known, deep learning performs well in a wide range of research and application areas. Exploitation of nonlinearities can produce highly accurate outputs. However, one of the problems with deep learning is that the number of parameters is enormous. In order to run such networks in environments with limited computational resources, model compression methods have been studied to reduce the number of model parameters. In fact, there are many studies that aim to compress neural networks [1]. Quantization and pruning are typical examples of model compression methods. Quantization reduces the data size of parameters by reducing the number of bits required to represent floating-point variables on a computer. Pruning reduces the number of connections between nodes; weights with small contributions are set to zero. These methods often require hardware support to maintain high accuracy while reducing memory consumption at runtime [2]. Other model compression methods in recent years include tensor decompositions with a low-rank approximation for the convolutional layer of a neural network [3, 4, 5].
@@
Some studies also combined neural networks with Koopman operator theory [12, 13, 14]; some works reported that time evolution predictors with the Koopman operator show better performance in time series prediction tasks than conventional neural networks such as recurrent neural networks [15]. Another way to apply Koopman operator theory to neural networks is to consider the learning process of weights as a dynamical system. For example, in [16], the Koopman operator is constructed from the early-stage learning trajectory of weights. Then, the use of the constructed Koopman operator in the subsequent optimization process significantly reduces the learning time. Reference [17] discussed the pruning methods in a unified manner with the aid of the Koopman operator theory. Reference [18] reported that good performance is achieved by applying the Koopman theory to the learning process of a network called the deep equilibrium model (DEQ).
Nowadays, many studies consider the internal process of neural networks as dynamical systems. For example, reference [19] considered ‘Residual Network’ as a discretization of ordinary differential equations by the Euler’s method and shows the improvement of the accuracy of classification problems with the aid of another discretization method. Reference [20] proposed a network called neural ODE, whose internal architecture is the ordinary differential equation; this fact yields straightforwardly to apply existing time-evolution solvers in the forward propagation stage. In [21], the recurrent neural network is regarded as a discretized ordinary differential equation. Hence, it is possible to apply the theory of stability of dynamical systems to the problems of gradient vanishing and explosion during training of the recurrent neural network. These discussions yielded a new structure of the recurrent neural network that avoids gradient vanishing and explosion; the proposed network showed better results than existing methods such as long short-term memory (LSTM) networks.
@@
In this study, we first discuss the effects of nonlinearity in trained neural networks. Of course, it is difficult to investigate the nonlinear effects, and the naive EDMD can deal with only a limited number of variables because of the curse of dimensionality; the memory size increases exponentially with the number of variables. Hence, we here employ the so-called tensor-train (TT) format [22, 23, 24], which expands the scope of application of DMD and EDMD (for example, see [25, 26]). After the preliminary investigation for nonlinearity, we consider neural networks for classification problems as dynamical systems and perform further analysis. Then, we discuss the ability of the Koopman operators for model compression.
The remaining part of the present paper is as follows. Section 2 describes the existing theory and methods necessary for the proposed method. In section 3, we propose a model compression framework with the aid of the Koopman operator. Preliminary numerical investigations using the TT format are also given. Compression with the proposed Koopman framework is discussed in section 4, in which we also yield numerical comparisons with the proposed method and compare them with existing methods. Section 5 gives concluding remarks.
@@ 2 Prerequisite
This cost function represents the mean squared error between the true values of the time-evolved dictionary functions and the corresponding predicted values on each snapshot pair. In (6), $\bm{\Phi}(\bm{x}_{i})$ represents that a data vector $\bm{x}_{i}$ is converted to a higher dimensional vector constructed by the dictionary functions. This operation is sometimes called “lifting”. Minimizing the cost function $\ell(K)$ is a linear least-squares problem whose solution is given by
@@
where ${}^{\dagger}$ means the pseudo-inverse matrix. $\bm{\Phi}(X)$ and $\bm{\Phi}(Y)\in\mathbb{R}^{M\times L}$ are the following matrices of the lifted data points:
@@
3 Model compression with Koopman operator and investigations of nonlinearity
In this section, we propose a Koopman-based model compression method for neural networks. After explaining the neural network architecture, we perform a preliminary numerical investigation in which the effects of nonlinearity are checked with the TT format.
@@ 3.1 Overview of the proposed method
As introduced in section 1, many studies consider the internal process of neural networks as a dynamical system. Although the dynamical system is nonlinear, we can apply the EDMD algorithm in principle. If the EDMD algorithm captures the essential part of the dynamical system, we can deal with this part with a linear transformation using a matrix, albeit in an approximate observable space. Therefore, we expect to compress the neural network model by using existing compression methods for the matrix.
Based on the above ideas, the proposed method uses roughly the following procedure for model compression.
@@
where $U\in\mathbb{R}^{L\times L}$ and $V\in\mathbb{R}^{D\times D}$ are orthogonal matrices, and $\Sigma\in\mathbb{R}^{L\times D}$ is a block matrix whose upper left part is a diagonal matrix of singular values $\sigma_{1},\ldots,\sigma_{r}>0\ (r=\mathrm{rank}\ A)$ and the rest parts are zero matrices. Then, we select $s$ singular values obtained by the singular value decomposition in descending order and extract the elements related to them from the matrices after singular value decomposition. If $s$ is sufficiently small, it is possible to approximate $A$ by a matrix with only a small number of elements.
@@ 3.2 Concrete example of the proposed method
First, we train a neural network that solves a classification problem in advance. The classification problem is to recognize a $28\times 28$ pixel handwritten digit image from the MNIST dataset as a numerical value. The neural network is a fully connected network whose activation function is the ReLU function $h(x)=\max\{0,x\}$. The number of nodes in each layer is $(784,20,20,20,20,20,10)$, where the first is the input layer, the last is the output one, and the rest correspond to multiple intermediate layers. The activation function is not applied to the first intermediate and output layers; all nonlinear processing is performed inside the other intermediate layers. The settings for training the neural network are shown in table 1.
@@
After training the neural network, we compute the Koopman matrix that mimics the internal processing of the network using EDMD. The EDMD algorithm requires snapshot pairs of the dynamic system. It is impossible to directly pair the inputs and outputs of the neural network of interest since the snapshot pairs must have the same dimension. Therefore, we consider the inputs and outputs of two layers in the intermediate layers as the snapshot pairs of the dynamical system; they have the same dimension when training data are forwardly propagated to the trained neural network. In the experiment, we select the first and last layers of the intermediate layers to generate snapshot pairs. Then, the Koopman matrix mimics all the nonlinear processing of the neural network. Figure 1 shows an overview of the network structure and application of EDMD.
Figure 1: Neural network, dynamical system, and the partial replacement with the Koopman matrix.
@@ 3.3 Preliminary investigation of nonlinearity with the TT format
Before considering model compression, we investigate the roles of nonlinearity in the trained neural network, that is, to what extent the neural network partially replaced by the Koopman matrix replicates the original behavior of the intermediate layers.
@@
From these preliminary numerical experiments with the TT format, we expect the possibility of compression of neural networks with the Koopman operator.
3.4 Practical dictionary for model compression
Although the TT format is a powerful tool for high-dimensional systems, it is still difficult to apply it naively for the learning process such as fine-tuning. In our naive code, a prediction per data takes about $0.7$ seconds for the TT format with $N_{\mathrm{max}}=3$. The TT format tells us that the limited nonlinearity could be enough. Hence, we use the more limited dictionary in the following discussion.
@@ 3.4 Practical dictionary for model compression
To confirm this in more detail, we investigate the classification accuracy of the network and the prediction error of the internal time evolution when the number of RBF dictionary functions gradually increased. Figure 4 shows the results; the classification accuracy of the network reached a plateau, while the prediction errors of the internal time evolution did not fully decrease. Therefore, as discussed above, there are cases in which the classification accuracy can be maintained even if there is a certain amount of prediction errors in the time evolution. This fact suggests that model compression could be possible by approximating the internal time evolution of the neural network.
4 Numerical experiments for model compression
In this section, we perform model compression on a neural network by the proposed method. Its performance is compared with that of pruning, one of the existing model compression methods.
@@ 4.1 Experimental details
4.1.1 Model compression by the proposed method
Model compression with the proposed method is performed on the same neural network in section 3. We here use the Gaussian RBF (and identity mapping) dictionary, which makes it easier to adjust the number of dictionary functions than the monomial dictionary. Parameters and selection of center points for the Gaussian RBF are the same as in section 3.
@@ 4.1.1 Model compression by the proposed method
4.1.2 Model compression by existing methods
For comparison, we use the conventional pruning method. Although the neural network is the same as in section 3, it is compressed by the pruning, as follows. Pruning is a method of reducing parameters by selecting unimportant weights and removing them. There are mainly two types of pruning: unstructured pruning and structured pruning. Unstructured pruning only sets some weights to 0, which does not reduce the memory used by the network to perform inference, but it is likely to have high accuracy. By contrast, in structured pruning, weights are removed in units of rows, columns, channels, and so on. Hence, we can reduce the size of the network structure and the memory size required to run the inference while it gives less accuracy than unstructured pruning.
@@ 4.1.2 Model compression by existing methods
4.1.3 Comparison of performance after fine tuning
Hence, we compare the performance of the proposed method with the fine-tuned conventional methods. The training procedure is the same as the first training. Only the difference is the number of epochs; the number is 5 in the fine-tuning and 14 in the original training. In the following experiments, we do not consider the center point $\bm{c}$ and the parameter $\varepsilon$ for a Gaussian RBF as training parameters.
@@ 4.2.1 Before fine-tuning
Figure 5 shows the accuracy of test data of the neural network with model compression in each setting: the number of dictionary functions and the number of ranks, which are shown in the vertical and horizontal axes, respectively. Note that we plot the results without singular value decomposition for the uppermost part of figure 5. When the number of ranks is 20, which equals the dimension of the state variable, the singular value decomposition only increases the number of matrix elements and does not change the calculation results. Hence, the results without performing singular value decomposition are shown. The blank space corresponds to the region where the total number of matrix elements in the three matrices of the singular value decomposition are greater than that of the case without singular value decomposition.
@@
Then, we conclude that the proposed compression method with the Koopman matrix works well.
4.2.2 After fine-tuning
Next, we show the experimental results for the case of fine-tuning. Figure 9 shows the accuracy of the fine-tuned compressed model by the proposed method. Figure 10 plots the accuracy versus compression ratio for the proposed and the conventional methods. Figure 11 shows the accuracy of the validation data after fine-tuning and the best settings adopted for each compression ratio, as in figure 8.
@@ 5 Conclusion
In this study, we proposed a new model compression method for neural networks by applying Koopman operator theory. We performed model compression on a network for recognition tasks of the handwritten numbers. The experimental results showed that the performance of the proposed method is comparable to that of the conventional method when the models are highly compressed. Note that the proposed method removes the activation function in the intermediate layers; only the Koopman matrix acts on the intermediate process. Even in this simple replacement, the network works well.
There is room for further study in the case of other network structures and different compression methods of matrices. In addition to model compression, we expect the Koopman operator to extract useful information from neural networks. Our approach is based on the theory of dynamical systems, and there are many studies on this topic. We must tackle high-dimensional problems that stem from large neural networks. We believe that the Koopman operator theory and the TT format used in this work could help future works that provide insights into the internal mechanisms of black-box neural networks.

### v2

# Extraction of nonlinearity in neural networks with Koopman operator

@@ Abstract
Nonlinearity plays a crucial role in deep neural networks. In this paper, we investigate the degree to which the nonlinearity of the neural network is essential. For this purpose, we employ the Koopman operator, extended dynamic mode decomposition, and the tensor-train format. The Koopman operator approach has been recently developed in physics and nonlinear sciences; the Koopman operator deals with the time evolution in the observable space instead of the state space. Since we can replace the nonlinearity in the state space with the linearity in the observable space, it is a hopeful candidate for understanding complex behavior in nonlinear systems. Here, we analyze learned neural networks for the classification problems. As a result, the replacement of the nonlinear middle layers with the Koopman matrix yields enough accuracy in numerical experiments. In addition, we confirm that the pruning of the Koopman matrix gives sufficient accuracy even at high compression ratios. These results indicate the possibility of extracting some features in the neural networks with the Koopman operator approach.
@@ 1 Introduction
As is well known, deep learning performs well in a wide range of research and application areas [1, 2, 3]. Exploitation of nonlinearities can produce highly accurate outputs. However, one of the problems with deep learning is that the number of parameters is enormous. In order to run such networks in environments with limited computational resources, model compression methods have been studied to reduce the number of model parameters. In fact, there are many studies that aim to compress neural networks [4, 5]. Quantization and pruning are typical examples of model compression methods. Quantization reduces the data size of parameters by reducing the number of bits required to represent floating-point variables on a computer. Pruning reduces the number of connections between nodes; weights with small contributions are set to zero. These methods often require hardware support to maintain high accuracy while reducing memory consumption at runtime [6]. There are other types of pruning, i.e., the spectral pruning [7, 8]. Other model compression methods in recent years include tensor decompositions with a low-rank approximation for the convolutional layer of a neural network [9, 10, 11].
@@
Some studies also combined neural networks with Koopman operator theory [18, 19, 20, 21]; the basic idea is to employ the neural networks to construct dictionaries for the Koopman operator. Actually, some works reported that time evolution predictors with the Koopman operator show better performance in time series prediction tasks than conventional neural networks such as recurrent neural networks [22]. Another way to apply Koopman operator theory to neural networks is to consider the learning process of weights as a dynamical system. For example, in [23], the Koopman operator is constructed from the early-stage learning trajectory of weights. Then, the use of the constructed Koopman operator in the subsequent optimization process significantly reduces the learning time. Reference [24] discussed the pruning methods in a unified manner with the aid of the Koopman operator theory. Reference [25] reported that good performance is achieved by applying the Koopman theory to the learning process of a network called the deep equilibrium model (DEQ).
The interpretability of neural networks is also a hot topic. There are some types of approach, such as a circuit approach[26] and a differential equation approach[27]. The connection of the neural network with the dynamical system stems the ResNet [28], and nowadays, many studies consider the internal process of neural networks as dynamical systems. For example, reference [29] considered the ResNet as a discretization of ordinary differential equations by the Euler’s method and shows the improvement of the accuracy of classification problems with the aid of another discretization method. Reference [30] proposed a network called neural ODE, whose internal architecture is the ordinary differential equation; this fact yields straightforwardly to apply existing time evolution solvers in the forward propagation stage. In [31], the recurrent neural network is regarded as a discretized ordinary differential equation. Hence, it is possible to apply the theory of stability of dynamical systems to the problems of gradient vanishing and explosion during training of the recurrent neural network. These discussions yielded a new structure of the recurrent neural network that avoids gradient vanishing and explosion; the proposed network showed better results than existing methods such as long short-term memory (LSTM) networks. Furthermore, the viewpoint of the dynamical systems, including partial differential equations, has also been studied; for example, see [32, 33, 34]. One would expect that discussions based on the dynamical systems will provide a better understanding of deep learning in the future.
@@
In this study, we discuss the effects of nonlinearity in trained neural networks. Of course, it is difficult to investigate the nonlinear effects, and the naive EDMD can deal with only a limited number of variables because of the curse of dimensionality; the memory size increases exponentially with the number of variables. Hence, we here employ the so-called tensor-train (TT) format [35, 36, 37], which expands the scope of application of DMD and EDMD (for example, see [38, 39]). After the preliminary investigation for nonlinearity, we consider neural networks for classification problems as discrete-time dynamical systems and perform further analysis. Then, we discuss the extent to which Koopman operators extract the essence of neural networks.
The remaining part of the present paper is as follows. Section 2 describes the existing theory and methods necessary for the proposed method. In section 3, we propose a model replacement framework with the aid of the Koopman operator. Preliminary numerical investigations using the TT format are also given. The main discussion with the proposed framework is given in section 4. Section 5 gives concluding remarks.
@@ 2 Prerequisite
This cost function represents the empirical mean squared error between the true values of the time-evolved dictionary functions and the corresponding predicted values on each snapshot pair. In (6), $\bm{\Phi}(\bm{x}_{i})$ represents that a data vector $\bm{x}_{i}$ is converted to a higher dimensional vector constructed by the dictionary functions. This operation is sometimes called “lifting”. Minimizing the cost function $\ell(K)$ is a linear least-squares problem whose solution is given by
@@
where means the pseudo-inverse matrix. $\bm{\Phi}(X)$ and $\bm{\Phi}(Y)\in\mathbb{R}^{M\times L}$ are the following matrices of the lifted data points:
@@
In the EDMD algorithm, the size of the dictionary should be large enough to approximate the observable functions correctly. Hence, $L\gg D$ in general.
3 Replacement with Koopman operator and investigations of nonlinearity
In this section, we propose a replacement of the intermediate layers in neural networks with the Koopman matrix. After explaining the neural network architecture, we perform a preliminary numerical investigation in which the effects of nonlinearity are checked with the TT format.
@@ 3.1 Overview of the proposed method
As introduced in section 1, many studies consider the internal process of neural networks as a dynamical system. Although the dynamical system is nonlinear, we can apply the EDMD algorithm in principle. If the EDMD algorithm captures the essential part of the dynamical system, we can deal with this part with a linear transformation using a matrix, albeit in an approximate observable space. Then, we consider the intermediate process in the neural networks as a discrete-time dynamical system and replace it with only a matrix. To investigate the possibility of extracting the essential features with the Koopman operator approach, we propose the following procedure:
@@
where $U\in\mathbb{R}^{L\times L}$ and $V\in\mathbb{R}^{D\times D}$ are orthogonal matrices, and $\Sigma\in\mathbb{R}^{L\times D}$ is a block matrix whose upper left part is a diagonal matrix of singular values $\sigma_{1},\ldots,\sigma_{r}>0\ (r=\mathrm{rank}\ A)$ and the rest parts are zero matrices. Then, we select $s$ singular values obtained by the singular value decomposition in descending order and extract the elements related to them from the matrices after singular value decomposition. If a sufficiently small $s$ yields good performance, only restricted features are enough to explain the information processing in the neural networks.
@@ 3.2 Concrete example of the proposed method
First, we train a neural network that solves a classification problem in advance. The classification problem is to recognize a $28\times 28$ pixel handwritten digit image from the MNIST dataset as a numerical value [40, 41]. The neural network is a fully connected network whose activation function is the ReLU function $h(x)=\max\{0,x\}$. The number of nodes in each layer is $(784,20,20,20,20,20,10)$, where the first is the input layer, the last is the output one, and the rest correspond to multiple intermediate layers. The activation function is not applied to the first intermediate and output layers; all nonlinear processing is performed inside the other intermediate layers. The settings for training the neural network are shown in table 1.
@@
After training the neural network, we compute the Koopman matrix that mimics the internal processing of the network using EDMD. The EDMD algorithm requires snapshot pairs of the dynamic system. It is impossible to directly pair the inputs and outputs of the neural network of interest since the snapshot pairs must have the same dimension. Therefore, we consider the inputs and outputs of two layers in the intermediate layers as the snapshot pairs of the discrete-time dynamical system; they have the same dimension when training data are forwardly propagated to the trained neural network. In the experiment, we select the first and last layers of the intermediate layers to generate snapshot pairs. Then, the Koopman matrix mimics all the nonlinear processing of the neural network. Figure 1 shows an overview of the network structure and application of EDMD.
Figure 1: Neural network, discrete-time dynamical system, and the partial replacement with the Koopman matrix.
@@ 3.3 Preliminary investigation of nonlinearity with the TT format
We investigate the roles of nonlinearity in the trained neural network, that is, to what extent the neural network partially replaced by the Koopman matrix replicates the original behavior of the intermediate layers.
@@
From these preliminary numerical experiments with the TT format, we expect that the Koopman operator extracts the essential features of neural networks well.
3.4 Practical dictionary for more detailed experiments
Although the TT format is a powerful tool for high-dimensional systems, it is still difficult to apply it naively. In our naive code, a prediction per data takes about $0.7$ seconds for the TT format with $N_{\mathrm{max}}=3$. The TT format tells us that the limited nonlinearity could be enough. Hence, we use the more limited dictionary in the following discussion.
@@ 3.4 Practical dictionary for more detailed experiments
To confirm this in more detail, we investigate the classification accuracy of the network and the prediction error of the internal time evolution when the number of RBF dictionary functions gradually increased. Figure 4 shows the results; the classification accuracy of the network reached a plateau, while the prediction errors of the internal time evolution did not fully decrease. Therefore, as discussed above, there are cases in which the classification accuracy can be maintained even if there is a certain amount of prediction errors in the time evolution.
4 How much does the Koopman matrix extract the essence?
The results in the previous section indicate that we can replace the intermediate layers with the Koopman matrix. Then, we check how much information is extracted by the Koopman matrix. In order to investigate it, we evaluate the performance by compressing the matrix with singular value decomposition. If we can achieve high compression without significant performance loss, it means that the Koopman matrix contains the essential information in a compact way. Then, we compare the compression ability of the Koopman matrix with the conventional pruning method.
@@ 4.1 Experimental details
In the following numerical experiments, the same neural network in section 3 is employed. We here use the Gaussian RBF (and identity mapping) dictionary, which makes it easier to adjust the number of dictionary functions than the monomial dictionary. Parameters and selection of center points for the Gaussian RBF are the same as in section 3.
@@
For comparison, we use the naive conventional pruning method. Although the neural network is the same as in section 3, it is compressed by the pruning, as follows. Pruning is a method of reducing parameters by selecting unimportant weights and removing them. There are mainly two types of pruning: unstructured pruning and structured pruning. Unstructured pruning only sets some weights to 0, which does not reduce the memory used by the network to perform inference, but it is likely to have high accuracy. By contrast, in structured pruning, weights are removed in units of rows, columns, channels, and so on. Hence, we can reduce the size of the network structure and the memory size required to run the inference while it gives less accuracy than unstructured pruning.
@@ 4.2 Experimental results
Figure 5 shows the accuracy of test data of the neural network with model compression in each setting: the number of dictionary functions and the number of ranks, which are shown in the vertical and horizontal axes, respectively. Note that we plot the results without singular value decomposition for the uppermost part of figure 5. When the number of ranks is 20, which equals the dimension of the state variable, the singular value decomposition only increases the number of matrix elements and does not change the calculation results. Hence, the results without performing singular value decomposition are shown for the 20-rank cases. The blank space corresponds to the region where the total number of matrix elements in the three matrices of the singular value decomposition are greater than that of the case without singular value decomposition.
@@
In addition, we perform additional numerical experiments for the Fashion MNIST dataset[44]. The numerical results show the same behavior; see the numerical results in Appendix B. These results indicate that the Koopman matrix can grasp the essence of the intermediate layers in the neural network in a compact way; a small part of the elements contains the crucial information of the intermediate layers.
The purpose of this paper is not to obtain the best pruning performance. However, some readers may be interested in the pruning performance. Then, we show some numerical results for fine-tuning cases in Appendix C.
@@ 5 Conclusion
In this study, we investigated the ability of the Koopman operator approach to investigate the neural networks. We performed the replacement of the intermediate layers on a network for recognition tasks with the Koopman matrix. The experimental results show that the Koopman matrix achieves good performance even in high compression cases. Hence, the essential parts of the intermediate layers are included in the Koopman matrix compactly.
There is room for further study in the case of other network structures and different compression methods of matrices. Of course, we expect the Koopman operator to extract useful information from neural networks, and further studies are hopeful. Our approach is based on the theory of dynamical systems, and there are many studies on this topic, as introduced in section 1. Then, it is hoped that the approach with the dynamical system will advance our understanding of neural networks from a physics perspective. We are performing some preliminary experiments in our ongoing research. The experiments indicate that we can replace some parts of the not-so-large neural networks with Koopman matrices, as shown in the current work. Then, we must tackle high-dimensional problems that stem from large neural networks in the future. We believe that the Koopman operator theory and the TT format used in this work could help future works that provide insights into the internal mechanisms of black-box neural networks.
@@ Appendix A Construction and usage of Koopman matrix in the TT format
Appendix B Experiments for the Fashion MNIST data set
Figure 11: Accuracy for various compressed models on the Fashion MNIST data set. The corresponding MNIST results are shown in figure 4.
We perform additional numerical experiments on a different dataset. Here, we apply the same procedure in section 4 on the Fashion MNIST dataset[44]. The Fashion MNIST dataset is known to be a more difficult dataset to classify than the MNIST dataset. Since the input data size of the Fashion MNIST is the same as the MNIST, it is possible to use the same network architecture. The experimental settings are exactly the same as in section 4, and the results are shown in figure 11. This figure corresponds to figure 4 on the MNIST dataset. We see that similar behavior on the MNIST dataset, which suggests the consequences do not depend on the dataset.
Appendix C Comparison of performance after fine tuning
We here compare the performance of the proposed method with the fine-tuned conventional methods. The training procedure is the same as the first training. Only the difference is the number of epochs; the number is 5 in the fine-tuning and 14 in the original training. In the following experiments, we do not consider the center point $\bm{c}$ and the parameter $\varepsilon$ for a Gaussian RBF as training parameters.
Figure 12 shows the accuracy of the fine-tuned compressed model by the proposed method. Figure 13 plots the accuracy versus compression ratio for the proposed and the conventional methods. Figure 14 shows the accuracy of the validation data after fine-tuning and the best settings adopted for each compression ratio, as in figure 8.